# Dynamical Tidal Response of Schwarzschild Black Holes

### v1

@@ 1 Introduction
In this paper we study subleading in frequency tidal effects, namely the dynamical response. Specifically, we aim to compute the induced time-dependent tidal deformation of Schwarzschild black holes by matching results obtained in general relativity to a point-particle EFT description of black holes [39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 6, 7, 13]. This provides an unambiguous definition of the dynamical Love numbers. The general topic of black hole dynamical response is well-studied [49, 50, 51, 32, 13, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69]. In particular, it is well known that frequency-dependent effects are nonzero. For example, a black hole absorbs radiation through its horizon, which, at linear order in frequency, corresponds to a positive imaginary part of the response coefficients [70, 13, 71]. Solutions to the perturbation equations at second order in frequency have also been obtained by various methods. These include the Mano–Suzuki–Takasugi (MST) formalism [72, 73]—which expresses the solution as a series of special functions truncated at a chosen order in the small-frequency limit—as well as more standard perturbative approaches [60, 67]. It is by now well understood that the dynamical Love numbers, unlike their static counterparts, exhibit a logarithmic dependence on distance, which is interpreted as a classical example of renormalization-group running in the point-particle EFT description of the object [49, 13, 56, 59, 60, 67, 74]. However, although the coefficient of the logarithm has been previously computed using the above methods, as well as other symmetry-based arguments [55], existing results in the literature remain incomplete for gravitational perturbations (see, however, [59, 74] for a scalar field tidal matching). In particular, a full computation of the (renormalized) dynamical Love numbers cannot be performed at tree level, but requires evaluating certain (classical) higher-loop diagrams in the point-particle EFT after subtracting ultraviolet divergences.
@@ 2.1 Small-frequency expansion of relativistic solution
We are therefore motivated to work in the adiabatic regime. That is, we will assume $\omega r_{s}\ll 1$ and seek solutions to (2.3) perturbatively in $\omega r_{s}$. It is important to note that, regardless of how small the frequency is, the term $r^{3}\omega^{2}/(r-r_{s})$ in (2.3) is not always guaranteed to be small. For instance, sufficiently close to the horizon ($r\rightarrow r_{s}$), this term dominates and eventually becomes the leading component of the potential. Likewise, this term grows at large distances, as $r\rightarrow\infty$. As such, the small frequency approximation is necessarily more complicated than simply neglecting the $r^{3}\omega^{2}/(r-r_{s})$ term. A natural way to proceed is to adopt an asymptotic expansion approach, reducing the problem to a simplified set of equations by making approximations which are valid in different regions with some overlap. The complete solution can then be reconstructed by matching the various solutions at the boundaries of their respective regions of validity. [^3] This type of approach is sometimes referred to as “boundary layer theory”. See [87] for a more rigorous introduction to the theory of boundary layer problems. See also [88, 89, 90, 91, 92] for a non-exhaustive list of scalar field applications on various black hole spacetimes.
@@
In each of these regions, we will be able to neglect certain terms in the original equation (2.3), allowing us to find an approximate, analytic closed-form solution. A central conceptual point of the analysis is that in order to match to point-particle EFT, it suffices to know the solution in the Intermediate Zone, as this has an overlapping regime of validity with the EFT description, if one matches the one-point function. In [74] they match to the EFT in the Far Zone, by matching scattering amplitudes to scattering computed in GR using the MST formalism. In order to make contact with this description, we show in Appendix B that the solution at small frequency reproduces the results of MST when matched to the Far Zone.
@@ 2.1.1 Near zone
Imposing ingpoing boundary conditions at the horizon, the physical near-zone solution thus reads
@@ 2.1.2 Intermediate zone
for some integration constants $b_{1}^{(k)}$ and $b_{2}^{(k)}$. [^5] The constants $b_{1}^{(k)}$ and $b_{2}^{(k)}$ are $\ell$-dependent, as we will explicitly see in some examples below. These can be determined by matching the perturbative solutions $\phi_{\ell}^{(0)}(x)$ and $\phi_{\ell}^{(1)}(x)$ to the near-zone solution (2.7) across the boundary at $r=r_{s}$. To this end, it is useful to recall the asymptotic behavior of the Legendre polynomials near $x=1$ (see e.g., [93, 94]):
@@
and where $\phi_{\ell,\text{p}}^{(2)}(x)$ is a particular solution to the equation, which can be expressed as a convolution between the Green’s function of the homogeneous equation and the source, as follows: [^6] We have used that the combination $W\equiv(1-x^{2})[P_{\ell}(x)\partial_{x}Q_{\ell}(x)-Q_{\ell}(x)\partial_{x}P_{\ell}(x)]$, proportional to the Wronskian of the homogeneous equation, is independent of $x$ for all $\ell$’s, i.e. $\partial_{x}W=0$. In particular, it is straightforward to check that $W\equiv 1$.
@@ 2.2.1 EFT setup
Here we have defined the traceless combination of derivatives $\nabla^{(\ell)}_{A_{\ell}}\equiv\nabla_{(i_{1}}\cdots\nabla_{i_{\ell})_{T}}$, where $(\cdots)_{T}$ denotes the symmetrized trace-free part of the enclosed indices. [^9] Note that (2.31) is written in terms of spatial indices, and so does not look covariant. This can be rectified by defining the projected covariant derivative $\nabla^{\bot}_{\mu}\equiv{\mathbb{P}^{\nu}}_{\mu}\nabla_{\nu}$, where ${\mathbb{P}^{\nu}}_{\mu}\equiv\delta_{\mu}^{\nu}+u_{\mu}u^{\nu}$ projects onto the plane orthogonal to the four velocity $u^{\mu}\equiv{\rm d}x^{\mu}/{\rm d}\tau$ (where $\tau$ is the proper time, and normalized such that $u^{\mu}u_{\mu}=-1$). The composite operator multipole operator $Q_{I}^{A_{\ell}}(\tau)$ is built from internal degrees of freedom $X_{\pm}$ localized on the worldline of the particle, which physically we can think of as modeling the degrees of freedom into which energy and other charges are dissipating. [^10] Introducing the auxiliary worldline degrees of freedom $X$ is not strictly necessary. One can instead directly write down the effective couplings (2.34) between the $\pm$ fields, following the Schwinger–Keldysh approach [107, 108, 109, 110, 111]. The two approaches are equivalent and lead to the same description of the open system. Correlation functions of the effective operator $Q$ describe how the point particle reacts to the presence of an external field $\Phi$.
@@
where we have taken advantage of the fact that in the case of interest the system is rotationally symmetric. [^11] This means that the Green’s function (2.34) is diagonal in $\ell,\ell^{\prime}$ and its index structure can be written in terms of $\delta_{ij}$. The end result is that the indices on the two derivative operators acting on $\Phi$ end up getting contracted. Correspondingly the responses at a given $\ell$ are independent of the magnetic quantum number $m$. The function $K^{(\Phi)}$ encodes the response of the particle to external sources. In frequency space, real terms capture conservative responses and imaginary terms capture dissipative response.
@@ 2.2.2 Linear response
Simplifying the ratio of gamma functions and writing the tidal field in spherical coordinates as $c_{i_{1}\cdots i_{\ell}}x^{i_{1}}\cdots x^{i_{\ell}}=c_{\text{ext}}r^{\ell}Y_{\ell}^{m}$, we obtain
@@
$\mkern 2.5mu\overline{\mkern-2.5mu\Phi\mkern-2.5mu}\mkern 2.5mu_{+}+\langle\Phi_{+}\rangle_{\text{in-in}}=c_{\text{ext}}\operatorname{e}^{-i\omega t}\sum_{\ell=0}^{\infty}Y_{\ell}^{m}\left(r^{\ell}+\mu^{2\varepsilon}K^{(\Phi)}_{\ell}(\omega)\frac{2^{\ell-2}\ell!\Gamma\left(\ell+\frac{D-3}{2}\right)}{\pi^{\frac{D-1}{2}}}r^{-\ell-D+3}\right).$ (2.45)
@@ 2.2.3 Coupling to gravity via the Born series
with $B_{\text{reg}}$ and $B_{\text{irr}}$ the two arbitrary integration constants. [^14] For convenience, we extracted the constant prefactor $\frac{\mu^{2\epsilon}}{2\ell+1-2\epsilon}$ from the decaying branch. Implicitly, the $r$-independent coefficients $B_{\text{reg}}$ and $B_{\text{irr}}$ should be thought of as expanded in powers of $G$ and $\omega$, as we discussed, and will see explicitly later on. The former branch of solution is regular everywhere in space, including the origin, with $B_{\text{reg}}$ corresponding to the amplitude of the asymptotic external tidal field that we use to probe the point object. On the other hand, the decaying $r^{-\ell+\varepsilon}$ solution is divergent at the location of the particle. We fix its free coefficient $B_{\text{irr}}$ by demanding that the divergence is sourced precisely by the localized action term $S_{\text{int}}$ (2.31). In practice, we determine the Love numbers $K_{\ell}^{(\Phi)}(\omega)$ by matching the homogeneous solution for $\Phi$
$\Phi_{+}^{(h)}=\operatorname{e}^{-i\omega t}\sum_{\ell=0}^{\infty}Y_{\ell}^{m}\,B_{\rm reg}\left(r^{\ell}+\frac{B_{\rm irr}}{B_{\rm reg}}\frac{\mu^{2\epsilon}}{2\ell+1-2\varepsilon}r^{-\ell-1+2\varepsilon}\right),$ (2.52)
@@ 3 Dynamical Love numbers
We first construct perturbative-in-frequency solutions to the Regge–Wheeler and Zerilli equations, paralleling the scalar analysis of Section 2 (see also [60, 67] for related treatments). We then match these solutions to the worldline EFT in dimensional regularization, and explicitly compute the dynamical Love number coefficients, including, for the first time, the finite scheme-dependent terms.
@@ 3.1 General relativistic solution via small-frequency expansion
We are interested in the dynamics of perturbations around a background Schwarzschild geometry (2.26). Denoting it with $\mkern 2.5mu\overline{\mkern-2.5mug\mkern-2.5mu}\mkern 2.5mu_{\mu\nu}$, we perturb the metric as ${g}_{\mu\nu}=\mkern 2.5mu\overline{\mkern-2.5mug\mkern-2.5mu}\mkern 2.5mu_{\mu\nu}+h_{\mu\nu}$. To make maximal use of the ${\rm SO}(D-1)$ background symmetry, it is convenient to decompose $h_{\mu\nu}$ into scalar, vector, and tensor spherical harmonics. Concretely, we can write $h_{\mu\nu}$ as follows [119, 120, 121, 7, 122]:
@@ 3.2.1 Odd sector
$=\frac{1}{\ell!}\int\frac{{\rm d}{\omega}}{2\pi}\operatorname{e}^{-i{\omega}(\tau_{2}-\tau_{1})}\left[\frac{\lambda_{0,\ell}^{B}}{2}+i{\omega}r_{s}\frac{\lambda_{1,\ell}^{B}}{2}+({\omega}r_{s})^{2}\frac{\lambda_{2,\ell}^{B}}{2}+\cdots\right]\,$,
where the parameters $\lambda_{i}$ are the response coefficients. Terms with odd powers of $\omega$ (which are time-reversal odd) capture the dissipative response of the object, while those with even powers of $\omega$ parametrize the conservative tidal deformability. In particular, $\lambda_{0,\ell}^{B}$ is related to the magnetic static Love numbers, while $\lambda_{2,\ell}^{B}$ corresponds to the quadratic-in-frequency dynamical Love numbers that we wish to compute.
To determine the one-point function (3.51), we need to compute the two-point function of $B$. We first express the magnetic component of the Weyl tensor in terms of the graviton field. Since the linearized Weyl tensor on Minkowski space is a gauge-invariant quantity, it is useful to fix a gauge. A convenient choice is the de Donder gauge, defined by $\partial^{\mu}(h_{\mu\nu}-\frac{1}{2}\eta_{\mu\nu}h)=0$, with $h\equiv\eta^{\mu\nu}h_{\mu\nu}$. In this gauge, the linearized equations of motion are $\square h_{\mu\nu}=0$, and the on-shell magnetic component of the Weyl tensor reads
$B_{abc}=C_{0abc}=2\left(\partial_{a}\partial_{[b}h_{c]0}-\partial_{0}\partial_{[b}h_{c]a}\right).$ (3.54)
@@
$B_{A_{\ell}\,j}=\partial_{i_{L}}h_{0j}-\partial_{j}\partial_{(i_{1}}\cdots\partial_{i_{\ell-1}}h_{i_{\ell})_{T}0}+\partial_{0}F_{A_{\ell}\,j}\,$, (3.55)
@@
${B}_{-\,A_{\ell}j}\mkern 2.5mu\overline{\mkern-2.5mu{B}\mkern-2.5mu}\mkern 2.5mu_{+}^{A_{\ell}\,j}=$   $2\frac{\ell+1}{\ell}\partial_{A_{\ell}}h_{0j}\partial^{A_{\ell-1}}\partial^{[i_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu^{j]}{}_{0}+\partial_{A_{\ell}}h_{0j}\partial_{0}\mkern 2.5mu\overline{\mkern-2.5muF\mkern-2.5mu}\mkern 2.5mu^{A_{\ell}\,j}-(\partial_{j}\partial_{(A_{\ell-1}}h_{i_{\ell})_{T}0})\partial_{0}\mkern 2.5mu\overline{\mkern-2.5muF\mkern-2.5mu}\mkern 2.5mu^{A_{\ell}\,j}$ (3.58)
$+\partial_{A_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{0j}\partial_{0}F^{A_{\ell}\,j}-(\partial_{j}\partial_{(A_{\ell-1}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{j_{\ell})_{T}0})\partial_{0}F^{A_{\ell}\,j}+O(\omega^{2}).$
@@
${B}_{-\,A_{\ell}\,j}\mkern 2.5mu\overline{\mkern-2.5mu{B}\mkern-2.5mu}\mkern 2.5mu_{+}^{A_{\ell}\,j}=$   $2\frac{\ell+1}{\ell}\partial_{A_{\ell}}h_{0j}\partial^{A_{\ell-1}}\partial^{[i_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu^{j]}{}_{0}$ (3.59)
@@
$\langle B_{+\,abc}{B}_{-\,A_{\ell}\,j}\rangle\mkern 2.5mu\overline{\mkern-2.5mu{B}\mkern-2.5mu}\mkern 2.5mu_{+}^{A_{\ell}\,j}=$ $\,\,\frac{4(\ell+1)}{\ell}\langle(\partial_{a}\partial_{[b}h_{+\,c]0})(\partial_{A_{\ell}}h_{-\,j0})\rangle\left[\partial^{A_{\ell-1}}\partial^{[i_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{+}^{j]}{}_{0}-\partial_{0}\partial^{A_{\ell-2}}\partial^{[i_{\ell-1}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{+}^{j]i_{\ell}}\right]$ (3.60)
$-\frac{4(\ell+1)}{\ell}\langle(\partial_{a}\partial_{[b}h_{+\,c]0})(\partial_{0}\partial_{A_{\ell-2}}\partial_{[i_{\ell-1}}h_{-\,j]i_{\ell}})\rangle\partial^{A_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{+}^{j}{}_{0}$
$-\frac{4(\ell+1)}{\ell}\langle(\partial_{0}\partial_{[b}h_{+\,c]a})(\partial_{A_{\ell}}h_{-\,0j})\rangle\partial^{A_{\ell-1}}\partial^{[i_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{+}^{j]}{}_{0}+O(\omega^{2})$

### v2

@@ 1 Introduction
In this paper we study subleading-in-frequency tidal effects, namely the dynamical response. Specifically, we aim to compute the induced time-dependent tidal deformation of Schwarzschild black holes by matching results obtained in general relativity to a point-particle EFT description of black holes [39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 6, 7, 13]. This provides an unambiguous definition of the dynamical Love numbers. The general topic of black hole dynamical response is well-studied [49, 50, 51, 32, 13, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69]. In particular, it is well known that frequency-dependent effects are nonzero. For example, a black hole absorbs radiation through its horizon, which, at linear order in frequency, corresponds to a positive imaginary part of the response coefficients [70, 13, 71]. Solutions to the perturbation equations at second order in frequency have also been obtained by various methods. These include the Mano–Suzuki–Takasugi (MST) formalism [72, 73]—which expresses the solution as a series of special functions truncated at a chosen order in the small-frequency limit—as well as more standard perturbative approaches [60, 67]. It is by now well understood that the dynamical Love numbers, unlike their static counterparts, exhibit a logarithmic dependence on distance, which is interpreted as a classical example of renormalization-group running in the point-particle EFT description of the object [49, 13, 56, 59, 60, 67, 74]. However, although the coefficient of the logarithm has been previously computed using the above methods, as well as other symmetry-based arguments [55], existing results in the literature remain incomplete for gravitational perturbations (see, however, [59, 74] for a scalar field tidal matching). In particular, a full computation of the (renormalized) dynamical Love numbers cannot be performed at tree level, but requires evaluating certain (classical) higher-loop diagrams in the point-particle EFT after subtracting ultraviolet divergences.
@@ 2.1 Small-frequency expansion of relativistic solution
We are therefore motivated to work in the adiabatic regime. That is, we will assume $\omega r_{s}\ll 1$ and seek solutions to (2.3) perturbatively in $\omega r_{s}$. It is important to note that, regardless of how small the frequency is, the term $r^{3}\omega^{2}/(r-r_{s})$ in (2.3) is not always guaranteed to be small. For instance, sufficiently close to the horizon ($r\rightarrow r_{s}$), this term dominates and eventually becomes the leading component of the potential. Likewise, this term grows at large distances, as $r\rightarrow\infty$. As such, the small frequency approximation is necessarily more complicated than simply neglecting the $r^{3}\omega^{2}/(r-r_{s})$ term. A natural way to proceed is to adopt an asymptotic expansion approach, reducing the problem to a simplified set of equations by making approximations which are valid in different regions. The complete solution can then be reconstructed by matching the various solutions at the boundaries of their respective regions of validity. [^3] This type of approach is sometimes referred to as “boundary layer theory”. See [87] for a more rigorous introduction to the theory of boundary layer problems. See also [88, 89, 90, 91, 92] for a non-exhaustive list of scalar field applications on various black hole spacetimes.
@@
In each of these regions, we will be able to neglect certain terms in the original equation (2.3), allowing us to find an approximate, analytic closed-form solution. A central conceptual point of the analysis is that in order to match to point-particle EFT, it suffices to know the solution in the Intermediate Zone, as this has an overlapping regime of validity with the EFT description, if one matches the one-point function. [^4] Note that what we are calling the “Intermediate Zone” is called the “Near Zone” in much of the literature. Perhaps what we call the Near Zone would better be called the “Very Near Zone”. We apologize for any confusion. In [74] they match to the EFT in the Far Zone, by matching scattering amplitudes to scattering computed in GR using the MST formalism. In order to make contact with this description, we show in Appendix B that the solution at small frequency reproduces the results of MST when matched to the Far Zone.
@@ 2.1.1 Near zone
Imposing ingoing boundary conditions at the horizon, the physical near-zone solution thus reads
@@ 2.1.2 Intermediate zone
for some integration constants $b_{1}^{(k)}$ and $b_{2}^{(k)}$. [^6] The constants $b_{1}^{(k)}$ and $b_{2}^{(k)}$ are $\ell$-dependent, as we will explicitly see in some examples below. These can be determined by matching the perturbative solutions $\phi_{\ell}^{(0)}(x)$ and $\phi_{\ell}^{(1)}(x)$ to the near-zone solution (2.7) across the boundary at $r=r_{s}$. [^7] As noted in [88], formally the approximations that we have made to define the Near and Intermediate zones have no region of overlap. Nevertheless, this does not preclude us from matching across their interface [87]. To this end, it is useful to recall the asymptotic behavior of the Legendre polynomials near $x=1$ (see e.g., [93, 94]):
@@
and where $\phi_{\ell,\text{p}}^{(2)}(x)$ is a particular solution to the equation, which can be expressed as a convolution between the Green’s function of the homogeneous equation and the source, as follows: [^8] We have used that the combination $W\equiv(1-x^{2})[P_{\ell}(x)\partial_{x}Q_{\ell}(x)-Q_{\ell}(x)\partial_{x}P_{\ell}(x)]$, proportional to the Wronskian of the homogeneous equation, is independent of $x$ for all $\ell$’s, i.e. $\partial_{x}W=0$. In particular, one can check that $W\equiv 1$.
@@ 2.2.1 EFT setup
Here we have defined the traceless combination of derivatives $\nabla^{(\ell)}_{A_{\ell}}\equiv\nabla_{(i_{1}}\cdots\nabla_{i_{\ell})_{T}}$, where $(\cdots)_{T}$ denotes the symmetrized trace-free part of the enclosed indices. [^11] Note that (2.31) is written in terms of spatial indices, and so does not look covariant. This can be rectified by defining the projected covariant derivative $\nabla^{\bot}_{\mu}\equiv{\mathbb{P}^{\nu}}_{\mu}\nabla_{\nu}$, where ${\mathbb{P}^{\nu}}_{\mu}\equiv\delta_{\mu}^{\nu}+u_{\mu}u^{\nu}$ projects onto the plane orthogonal to the four velocity $u^{\mu}\equiv{\rm d}x^{\mu}/{\rm d}\tau$ (where $\tau$ is the proper time, and normalized such that $u^{\mu}u_{\mu}=-1$). The composite multipole operator $Q_{I}^{A_{\ell}}(\tau)$ is built from internal degrees of freedom $X_{\pm}$ localized on the worldline of the particle, which physically we can think of as modeling the degrees of freedom into which energy and other charges are dissipating. [^12] Introducing the auxiliary worldline degrees of freedom $X$ is not strictly necessary. One can instead directly write down the effective couplings (2.34) between the $\pm$ fields, following the Schwinger–Keldysh approach [107, 108, 109, 110, 111]. The two approaches are equivalent and lead to the same description of the open system. Correlation functions of the effective operator $Q$ describe how the point particle reacts to the presence of an external field $\Phi$.
@@
where we have used the fact that in the case of interest the system is rotationally symmetric. [^13] This means that the Green’s function (2.34) is diagonal in $\ell,\ell^{\prime}$ and its index structure can be written in terms of $\delta_{ij}$. The end result is that the indices on the two derivative operators acting on $\Phi$ end up getting contracted. Correspondingly the responses at a given $\ell$ are independent of the magnetic quantum number $m$. The function $K^{(\Phi)}$ encodes the response of the particle to external sources. In frequency space, real terms capture conservative responses and imaginary terms capture dissipative response.
@@ 2.2.2 Linear response
Simplifying the ratio of gamma functions and writing the tidal field in spherical coordinates as $c_{i_{1}\cdots i_{\ell}}x^{i_{1}}\cdots x^{i_{\ell}}=c_{\text{ext}}r^{\ell}Y_{\ell}^{m}$, we obtain [^15] We suppress indices on $c_{\text{ext}}$ which depends implicitly on the spherical harmonic quantum numbers $\ell,m$.
@@
$\mkern 2.5mu\overline{\mkern-2.5mu\Phi\mkern-2.5mu}\mkern 2.5mu_{+}+\langle\Phi_{+}\rangle_{\text{in-in}}=\operatorname{e}^{-i\omega t}\sum_{\ell=0}^{\infty}c_{\text{ext}}Y_{\ell}^{m}\left(r^{\ell}+\mu^{2\varepsilon}K^{(\Phi)}_{\ell}(\omega)\frac{2^{\ell-2}\ell!\Gamma\left(\ell+\frac{D-3}{2}\right)}{\pi^{\frac{D-1}{2}}}r^{-\ell-D+3}\right).$ (2.45)
@@ 2.2.3 Coupling to gravity via the Born series
with $B_{\text{reg}}$ and $B_{\text{irr}}$ the two arbitrary integration constants. [^17] For convenience, we extracted the constant prefactor $\frac{\mu^{2\varepsilon}}{2\ell+1-2\varepsilon}$ from the decaying branch. Implicitly, the $r$-independent coefficients $B_{\text{reg}}$ and $B_{\text{irr}}$ should be thought of as expanded in powers of $G$ and $\omega$, as we discussed, and will see explicitly later on. The former branch of solution is regular everywhere in space, including the origin, with $B_{\text{reg}}$ corresponding to the amplitude of the asymptotic external tidal field that we use to probe the point object. On the other hand, the decaying $r^{-\ell+\varepsilon}$ solution is divergent at the location of the particle. We fix its free coefficient $B_{\text{irr}}$ by demanding that the divergence is sourced precisely by the localized action term $S_{\text{int}}$ (2.31). In practice, we determine the Love numbers $K_{\ell}^{(\Phi)}(\omega)$ by matching the homogeneous solution for $\Phi$
$\Phi_{+}^{(h)}=\operatorname{e}^{-i\omega t}\sum_{\ell=0}^{\infty}Y_{\ell}^{m}\,B_{\rm reg}\left(r^{\ell}+\frac{B_{\rm irr}}{B_{\rm reg}}\frac{\mu^{2\varepsilon}}{2\ell+1-2\varepsilon}r^{-\ell-1+2\varepsilon}\right),$ (2.52)
@@ 3 Dynamical Love numbers
We first construct perturbative-in-frequency solutions to the Regge–Wheeler and Zerilli equations, paralleling the scalar analysis of Section 2 (see also [60, 67] for related treatments). We then match these solutions to the worldline EFT in dimensional regularization, which allows us to explicitly compute the dynamical Love number coefficients, including, for the first time, the finite scheme-dependent terms.
@@ 3.1 General relativistic solution via small-frequency expansion
We are interested in the dynamics of perturbations around a background Schwarzschild geometry (2.26). Denoting it with $\mkern 2.5mu\overline{\mkern-2.5mug\mkern-2.5mu}\mkern 2.5mu_{\mu\nu}$, we perturb the metric as ${g}_{\mu\nu}=\mkern 2.5mu\overline{\mkern-2.5mug\mkern-2.5mu}\mkern 2.5mu_{\mu\nu}+2h_{\mu\nu}/M_{\rm Pl}$. To make maximal use of the ${\rm SO}(D-1)$ background symmetry, it is convenient to decompose $h_{\mu\nu}$ into scalar, vector, and tensor spherical harmonics. Concretely, we can write $h_{\mu\nu}$ as follows [119, 120, 121, 7, 122]:
@@ 3.2.1 Odd sector
$=\frac{1}{\ell!}\frac{\ell}{\ell+1}\int\frac{{\rm d}{\omega}}{2\pi}\operatorname{e}^{-i{\omega}(\tau_{2}-\tau_{1})}\left[\lambda_{0,\ell}^{B}+i{\omega}r_{s}\lambda_{1,\ell}^{B}+({\omega}r_{s})^{2}\lambda_{2,\ell}^{B}+\cdots\right]\,$,
where the parameters $\lambda_{i}$ are the response coefficients and the overall normalization is chosen for later convenience. Terms with odd powers of $\omega$ (which are time-reversal odd) capture the dissipative response of the object, while those with even powers of $\omega$ parametrize the conservative tidal deformability. In particular, $\lambda_{0,\ell}^{B}$ is related to the magnetic static Love numbers, while $\lambda_{2,\ell}^{B}$ corresponds to the quadratic-in-frequency dynamical Love numbers that we wish to compute.
To determine the one-point function (3.51), we need to compute the two-point function of $B$. We first express the magnetic component of the Weyl tensor in terms of the graviton field. Since the linearized Weyl tensor on Minkowski space is a gauge-invariant quantity, it is useful to fix a gauge. A convenient choice is the de Donder gauge, defined by $\partial^{\mu}(h_{\mu\nu}-\frac{1}{2}\eta_{\mu\nu}h)=0$, with $h\equiv\eta^{\mu\nu}h_{\mu\nu}$. In this gauge, the linearized equations of motion are $\square h_{\mu\nu}=0$, and the on-shell magnetic component of the Weyl tensor reads [^28] Recall that $h_{\mu\nu}$ is the canonically normalized graviton field.
$B_{abc}=C_{0abc}=\frac{2}{M_{\rm Pl}}\left(\partial_{a}\partial_{[b}h_{c]0}-\partial_{0}\partial_{[b}h_{c]a}\right).$ (3.54)
@@
$B_{A_{\ell}\,j}=\frac{1}{M_{\rm Pl}}\left[\partial_{i_{L}}h_{0j}-\partial_{j}\partial_{(i_{1}}\cdots\partial_{i_{\ell-1}}h_{i_{\ell})_{T}0}+\partial_{0}F_{A_{\ell}\,j}\right]\,$, (3.55)
@@
${B}_{-\,A_{\ell}j}\mkern 2.5mu\overline{\mkern-2.5mu{B}\mkern-2.5mu}\mkern 2.5mu_{+}^{A_{\ell}\,j}=\frac{1}{M_{\rm Pl}^{2}}\bigg[$ $2\frac{\ell+1}{\ell}\partial_{A_{\ell}}h_{0j}\partial^{A_{\ell-1}}\partial^{[i_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu^{j]}{}_{0}+\partial_{A_{\ell}}h_{0j}\partial_{0}\mkern 2.5mu\overline{\mkern-2.5muF\mkern-2.5mu}\mkern 2.5mu^{A_{\ell}\,j}-(\partial_{j}\partial_{(A_{\ell-1}}h_{i_{\ell})_{T}0})\partial_{0}\mkern 2.5mu\overline{\mkern-2.5muF\mkern-2.5mu}\mkern 2.5mu^{A_{\ell}\,j}$ (3.58)
$+\partial_{A_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{0j}\partial_{0}F^{A_{\ell}\,j}-(\partial_{j}\partial_{(A_{\ell-1}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{j_{\ell})_{T}0})\partial_{0}F^{A_{\ell}\,j}+O(\omega^{2})\bigg].$
@@
${B}_{-\,A_{\ell}\,j}\mkern 2.5mu\overline{\mkern-2.5mu{B}\mkern-2.5mu}\mkern 2.5mu_{+}^{A_{\ell}\,j}=\frac{1}{M_{\rm Pl}^{2}}\bigg[$ $2\frac{\ell+1}{\ell}\partial_{A_{\ell}}h_{0j}\partial^{A_{\ell-1}}\partial^{[i_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu^{j]}{}_{0}$ (3.59)
@@
$\langle B_{+\,abc}{B}_{-\,A_{\ell}\,j}\rangle\mkern 2.5mu\overline{\mkern-2.5mu{B}\mkern-2.5mu}\mkern 2.5mu_{+}^{A_{\ell}\,j}=$ $\,\,\frac{4(\ell+1)}{\ell M_{\rm Pl}^{3}}\langle(\partial_{a}\partial_{[b}h_{+\,c]0})(\partial_{A_{\ell}}h_{-\,j0})\rangle\left[\partial^{A_{\ell-1}}\partial^{[i_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{+}^{j]}{}_{0}-\partial_{0}\partial^{A_{\ell-2}}\partial^{[i_{\ell-1}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{+}^{j]i_{\ell}}\right]$ (3.60)
$-\frac{4(\ell+1)}{\ell M_{\rm Pl}^{3}}\langle(\partial_{a}\partial_{[b}h_{+\,c]0})(\partial_{0}\partial_{A_{\ell-2}}\partial_{[i_{\ell-1}}h_{-\,j]i_{\ell}})\rangle\partial^{A_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{+}^{j}{}_{0}$
$-\frac{4(\ell+1)}{\ell M_{\rm Pl}^{3}}\langle(\partial_{0}\partial_{[b}h_{+\,c]a})(\partial_{A_{\ell}}h_{-\,0j})\rangle\partial^{A_{\ell-1}}\partial^{[i_{\ell}}\mkern 2.5mu\overline{\mkern-2.5muh\mkern-2.5mu}\mkern 2.5mu_{+}^{j]}{}_{0}+O(\omega^{2})$